# Toward Tiny and High-quality Facial Makeup with Data Amplify Learning

### v1

@@ Abstract
Contemporary makeup approaches primarily hinge on unpaired learning paradigms, yet they grapple with the challenges of inaccurate supervision (e.g., face misalignment) and sophisticated facial prompts (including face parsing, and landmark detection). These challenges prohibit low-cost deployment of facial makeup models, especially on mobile devices. To solve above problems, we propose a brand-new learning paradigm, termed "Data Amplify Learning (DAL)," alongside a compact makeup model named "TinyBeauty." The core idea of DAL lies in employing a Diffusion-based Data Amplifier (DDA) to "amplify" limited images for the model training, thereby enabling accurate pixel-to-pixel supervision with merely a handful of annotations. Two pivotal innovations in DDA facilitate the above training approach: (1) A Residual Diffusion Model (RDM) is designed to generate high-fidelity detail and circumvent the detail vanishing problem in the vanilla diffusion models; (2) A Fine-Grained Makeup Module (FGMM) is proposed to achieve precise makeup control and combination while retaining face identity. Coupled with DAL, TinyBeauty necessitates merely 80K parameters to achieve a state-of-the-art performance without intricate face prompts. Meanwhile, TinyBeauty achieves a remarkable inference speed of up to 460 fps on the iPhone 13. Extensive experiments show that DAL can produce highly competitive makeup models using only 5 image pairs.
@@ 1 Introduction
The root cause behind the excessive model sizes and complex inference pipelines in current mainstream makeup models can be attributed to the underlying learning framework, which is inherently flawed. The annotation of paired makeup data is expensive and challenging to achieve style scalability. Consequently, unpaired data has become the mainstream data protocol, which requires makeup models to carefully employ adversarial training [18, 2, 14] for achieving stable unpaired learning. Moreover, unpaired data often suffers from significant facial misalignment issues, which necessitates the inclusion of facial spatial prompts and other operations (e.g., warping) to assist in the unpaired learning process. For example, explicit incorporation of facial landmarks is used to guide the network in identifying the position of lipstick, eyelashes, and eyeshadow. However, these additional prompts further contribute to the complexity of the model. Simultaneously, the data misalignment poses a challenge for unpaired learning models to employ stable and straightforward reconstruction loss (e.g., L1, L2). Instead, they rely on inaccurate supervision, such as color histograms matching [18], earth mover’s distance, due to the lack of alignment information. As well known, these inaccurate supervision methods often rely on theoretical approximations (e.g., approximate upper and lower bounds), leading to poor model robustness. For instance, as shown in Fig. Fig. 2, even with a model size of 10M, EleGANt [35] still frequently produces unsatisfactory results. The above challenges become more pronounced as the model size continues to decrease. The reduced model scale makes the model more sensitive to the accuracy of supervision while also requiring the abandonment of various auxiliary alignment techniques.
@@
To address the above challenges, we propose a novel learning framework named Data Amplify Learning enabling stable and robust training supervision even with a limited number of image annotations. In the heart of DAL is a meticulously designed Diffusion-based Data Amplifier, responsible for amplifying the limited labeled image pairs (typically 5 pairs) into a larger dataset suitable for model training. This process allows DAL to transform the inaccurate supervision into pixel-to-pixel learning (e.g., L1), reducing the optimization difficulty and enabling accurate gradient propagation. Naively using mainstream diffusion-based control methods [36, 30] as the Data Amplifier can lead to subpar outcomes, including identity mismatching, over-smoothing details, and inaccurate makeup results. To alleviate above challenges, two ingredients of the DDA are proposed: (1) Instead of directly generating makeup results, we design a Residual Diffusion Model (RDM) for the Data Amplifier. RDM consists of two branches: one is responsible for face reconstruction, and the other is dedicated to applying makeup. By utilizing the residual between the reconstructed face and the input face, we can transfer this residual to the makeup branch, resulting in high-fidelity and sharpened makeup results. (2) Unlike direct image injection [36], we propose a Fine-Grained Makeup Module (FGMM) that couples fine-grained makeup with facial identity representation to control makeup process, which effectively avoids the loss of facial identity. Additionally, we introduce semantic region labels for makeup, further refining the control of makeup application. rendering a precise makeup control generation. DAL greatly relaxes the optimization methods, allowing us to abandon face prompts and over-parameterization methods, thereby focusing on network design for mobile computing. Benefiting from this, a single TinyBeauty model is designed without reliance on external face prompts. TinyBeauty exceeds prior makeup transfer performance using only 14 convolution layers, enabling efficient deployment on mobile devices. Furthermore, to produce more complete makeup components compared to prior works, we utilize an edge operator to constrain the learning of eyeliner, enhancing the realism and fidelity of generated makeup styles.
@@ 2.1 Diffusion-based Image-to-Image Translation
Compared to previously widely used GAN-based [8] models like Pix2Pix [13] and CycleGAN [41], recent diffusion models [29, 5, 26] have demonstrated superior performance in high-fidelity image generation tasks, especially when few-shot images are available. Typically, conditions are introduced to steer the image generation process, generally in two forms: text-based and image-based guidance. DALL·E2 [20] and Imagen [27] apply text encoders from pre-trained language models [22, 4] for image generation guidance. Textual Inversion[7] and DreamBooth[25] learn special tokens from example images. However, solely using text for conditional guidance is indirect and inaccurate, making it difficult to ensure stable, consistently generated images. To strengthen conditional constraints, SD Image Variations2 [28] and Stable unCLIP3 [31] directly fine-tune text-conditioned diffusion models on image embeddings. IP-Adapter [36] and InstantID [33] use a combined image and text prompt for data generation conditions to keep the identity of the condition image. While these methods yield commendable outcomes in image stylization, they fail to preserve the facial detail like wrinkles of the original image, thus inadequate in applying facial makeup.
@@
$*{}^{*}$he latency is the inference time on an iPhone 13 device.
@@ 2.2 Facial Makeup
Traditional facial makeup methods [9, 17] rely on facial landmarks to warp predetermined beauty materials on the facial images for makeup application, which is efficient but unrealistic. Inspired by CycleGAN [41], BeautyGAN [18] introduces a dual GAN approach for makeup transfer with a color histogram loss using the proposed dataset. PSGAN [14] and FAT [32] propose attention mechanisms to address pose/expression changes, significantly increasing model size. SCGAN [2] encodes styles into component-wise codes while EleGANt [35] simplifies this complex optimization problem as L1 loss by generating pseudo ground truth. However, EleGANt still consumes significant computational overhead due to its use of multi-scale attention modules. By abandoning CycleGAN’s structure, BeautyREC [34] achieves a lighter model compared to previous work. Typically, all these methods incorporate parsing maps and/or landmarks during pre-processing on the unpaired MT dataset, forming cumbersome pipelines.
@@ 3 Method
We propose a novel learning scheme called Data Amplify Learning (DAL) to replace the unstable and inaccurate unpaired learning. DAL leverages a diffusion model to generate previously inaccessible paired data, which is then employed as training material for a compact tiny model. The core of DAL is a Diffusion-based Data Amplifier (DDA) to generate large amounts of paired data using only 5 makeup images as data seeds. (3.1). Benefit from the DDA-amplified data, a single TinyBeauty Model is designed to replace the previous cumbersome makeup pipeline, facilitating the integration of the makeup model into mobile devices. (3.2).
@@ 3.1 Diffusion-based Data Amplifier
$\vspace{-0.05cm}L_{simple}=\mathbb{E}_{\mathbf{z}_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},t}||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c},t)||^{2},$ (1)
@@
$\vspace{-0.05cm}y=\mathcal{F}_{fine}(x,\mathbf{c}),$ (2)
@@
3.1.1 Residual Diffusion Models
To address the issue of facial features like wrinkles being smoothed out during the denoising step of diffusion-based portrait image generation, we introduce a novel framework that utilizes parallel diffusion branches during the inference phase—the conditioned and unconditioned branches. The conditioned branch operates with conditions $\mathbf{c}_{con}$ and $\mathbf{c}_{sty}$ to produce a smooth makeup image, while the unconditioned branch proceeds without any conditions to output a smooth non-makeup image.
@@ 3.1.1 Residual Diffusion Models
3.1.2 Fine-Grained Makeup Module
The Fine-Grained Makeup Module comprises the Identity Preservation Block (IPB), which safeguards individual facial identity, the Style Preservation Block (SPB) for precise makeup style management, and a latent face mask that guides region-specific makeup application and facilitates the flawless integration of diverse makeup styles.
@@ 3.1.2 Fine-Grained Makeup Module
Mask Guidance. To accurately define the facial areas affected by the makeup style, we divide the feature space into three distinct regions under the guidance of resized facial masks: $M_{face}$, $M_{lips}$, and $M_{eyes}$, as shown in Fig. 4. For single makeup style learning, we define the influenced area $M_{changed}$ as $M_{changed}=M_{face}+M_{lips}+M_{eyes}$. In the training process, we only focus the learning on $M_{changed}$ regions to effectively capture the makeup style. This is achieved by computing the loss $L_{simple}$ only over the features within the $M_{changed}$ area:
$\vspace{-0.05cm}L_{simple}^{M}=\mathbb{E}_{\mathbf{z}_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},t}||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c},t)||^{2}*M_{changed}.\vspace{-0.05cm}$ (4)
@@ 3.2 TinyBeauty Model
$\vspace{-0.05cm}\mathcal{L}_{s}=||\mathcal{S}(y)-\mathcal{S}(y^{\prime})||_{2}^{2}*M_{eyes},\vspace{-0.05cm}$ (5)
@@
$\vspace{-0.05cm}\mathcal{L}_{rec}=||y-y^{\prime}||_{1}.\vspace{-0.05cm}$ (6)
@@ 4.1 Datasets and Evaluation Metrics
Data Seeds and Evaluation Dataset. We randomly select 5 images without makeup from the Flickr-Faces-HQ (FFHQ) [15] dataset as our DAL’s data seeds, i.e., training data. We then manually annotate these data seeds with five predefined makeup styles using MEITU  [^1] https://www.meitu.com/, along with the makeup-only images as the input of SPB. To enable a comprehensive and fair evaluation, 100 images from the FFHQ dataset [15] and 100 images from the MT dataset [18] are additionally annotated as test data when computing PSNR, FID, and LPIPS.
@@
Evaluation Metrics. To evaluate TinyBeauty quantitatively, we employ widely recognized metrics to evaluate the similarity, diversity, and realism of the images generated. We use the Peak Signal-to-Noise Ratio (PSNR) to measure the similarity between the generated data and ground truth (GT) data. A higher PSNR indicates a closer approximation to the GT. Additionally, we report the Fréchet Inception Distances (FID)  [10] and Learned Perceptual Image Patch Similarity (LPIPS) [39] by comparing 100 generated images to 100 GT images. FID and LPIPS metrics evaluate the diversity and realism of the generated samples, with lower scores indicating a higher similarity to the GT data distribution.
@@ 4.2 Implementation Details
Diffusion-based Data Amplifier (DDA). Our diffusion fine-tuning experiment use SDv1.5 [21] as the base model. The image encoder utilized is OpenCLIP ViT-H/14 [6], which is pre-trained by IP-Adapter [36]. We set the length of both style token and identity token to 32 to balance the strength of style and identity. LoRA [12] in conjunction with U-net is employed to fine-tune the SD model with a learning rate of 1e-4 for 500 epochs. Notably, the five distinct makeup styles are concurrently trained within a unified model, demonstrating the model’s capacity for multi-style learning in a single training session. The training process is executed on an NVIDIA V100 GPU, taking about 50 mins to fine-tune five makeup styles. We utilize the FaRL [40] to generate facial masks and resize them to a resolution of 64x64 to guide the network’s training in the latent space.
TinyBeauty Model. The TinyBeauty model is trained using a dataset consisting of 4000 diffusion-generated images for 50 epochs. The training processes utilized a learning rate of 2e-4 and the Adam [16] optimizer. The entire training procedure is executed on an NVIDIA V100 GPU, requiring approximately 12 hours. It is important to mention that NO specific face operations are conducted on the input image during the inference phase, which makes our makeup process fast and hardware-friendly.
@@ 4.3 Comparison
To showcase the quality of the data generated using our DDA, we compare it with the text-conditioned SD model, ControlNet [38], IP-Adapter [36], and InstantID [33]. We train both previous methods in the same data as ours, with the result shown in Fig. 5. The results reveal two key insights. Firstly, the subtleties of makeup styles are not easily captured by textual descriptions alone. Consequently, it is challenging to precisely manipulate attributes, particularly color, in the synthesized images. Our DDA addresses this by directly encoding makeup images, thus producing results with consistent color and makeup content. Secondly, IP-Adapter falls short of meeting our stringent criteria for identity retention in the facial makeup task. Our DDA, however, effectively resolves this issue, generating high-quality makeup images by synergistically leveraging mask guidance, IPB, and RDM. Additionally, 1 illustrates the DDA’s capability of fusing various makeup styles. Our DDA adeptly blends eye, lip, and blush makeup to create a myriad of diverse and intricate makeup applications.
@@
Table 2: Parameters of TinyBeauty and competing methods. (+) means the method uses facial pre-processing including face $*{}^{*}$, and (-) means the method only has a single model.
@@
Quantitative Comparison. We conduct a comprehensive evaluation by comparing TinyBeauty not only with the pre-trained models provided by earlier studies [18, 34, 14, 2, 35], but also by retraining the two most advanced models [35, 34] from those studies using our DAL. The comparative analysis, detailed in 1, reveals TinyBeauty’s performance over the pre-established and newly trained models in terms of PSNR, FID, and LPIPS across both the FFHQ and MT datasets. Remarkably, despite its significantly smaller network size, TinyBeauty achieves comparable results to EleGANt in the same training set. The success lies in the high-quality paired data we generate, which enables even small networks to learn and apply stable makeup styles effectively. Moreover, Our specialized design incorporating residual learning not only diminishes noise interference but also enhances TinyBeauty’s performance, even surpassing larger predecessors. Notably, the MT dataset is not utilized during training, yet TinyBeauty yields a PSNR improvement of 1.49dB over prior methods on MT.
Model Efficiency Evaluation. As shown in Fig. 2, we measure FLOPs, parameter size, and inference time of each model. The models are packaged with CoreML [1] and tested on an iPhone13 without including face pre-processing time. We also test a face parsing model [37] on iPhone13 for reference, commonly utilized by other approaches but not integrated within our TinyBeauty method. Our model achieves an inference time of only 2.18ms for a 256x256 image, which is 13×13$\times$13 × faster than the fastest competing method on iPhone13, i.e., BeautyGAN [18]. Specifically, TinyBeauty is approximately 6 times faster than face parsing on iPhone hardware, which is used as a pre-processing step in previous makeup transfer pipelines [34, 2, 35, 14]. In contrast, our TinyBeauty only utilizes a single model by simplifying the optimization problem facilitated by amplified paired data. These results highlight TinyBeauty’s efficiency and mobile deployment readiness, achieved by implementing our DAL scheme.
@@
User Study. We conduct a subjective evaluation on the FFHQ and MT test sets. Workers are shown makeup results from TinyBeauty and competing methods for raw images in random order. They are asked to rank the quality based on accurate makeup color, clear details like eyebrows or eyeliner, pleasing effects, and minimal distortion. A total of 100 validated workers participate in the fair and randomized evaluation. The average rankings are summarized in Fig. 3.

### v4

@@ Abstract
Contemporary makeup approaches primarily hinge on unpaired learning paradigms, yet they grapple with the challenges of inaccurate supervision (e.g., face misalignment) and sophisticated facial prompts (including face parsing, and landmark detection). These challenges prohibit low-cost deployment of facial makeup models, especially on mobile devices. To solve above problems, we propose a brand-new learning paradigm, termed "Data Amplify Learning (DAL)," alongside a compact makeup model named "TinyBeauty." The core idea of DAL lies in employing a Diffusion-based Data Amplifier (DDA) to "amplify" limited images for the model training, thereby enabling accurate pixel-to-pixel supervision with merely a handful of annotations. Two pivotal innovations in DDA facilitate the above training approach: (1) A Residual Diffusion Model (RDM) is designed to generate high-fidelity detail and circumvent the detail vanishing problem in the vanilla diffusion models; (2) A Fine-Grained Makeup Module (FGMM) is proposed to achieve precise makeup control and combination while retaining face identity. Coupled with DAL, TinyBeauty necessitates merely 80K parameters to achieve a state-of-the-art performance without intricate face prompts. Meanwhile, TinyBeauty achieves a remarkable inference speed of up to 460 fps on the iPhone 13. Extensive experiments show that DAL can produce highly competitive makeup models using only 5 image pairs. Please visit https://github.com/TinyBeauty for code and demos.
@@ 1 Introduction
The root cause behind the excessive model sizes and complex inference pipelines in current mainstream makeup models can be attributed to the underlying learning framework, which is inherently flawed. The annotation of paired makeup data is expensive and challenging to achieve style scalability. Consequently, unpaired data has become the mainstream data protocol, which requires makeup models to carefully employ adversarial training [21, 6, 17] for achieving stable unpaired learning. Moreover, unpaired data often suffers from significant facial misalignment issues, which necessitates the inclusion of facial spatial prompts and other operations (e.g., warping) to assist in the unpaired learning process. For example, explicit incorporation of facial landmarks is used to guide the network in identifying the position of lipstick, eyelashes, and eyeshadow. However, these additional prompts further contribute to the complexity of the model. Simultaneously, the data misalignment poses a challenge for unpaired learning models to employ stable and straightforward reconstruction loss (e.g., L1, L2). Instead, they rely on inaccurate supervision, such as color histograms matching [21], earth mover’s distance, due to the lack of alignment information. As well known, these inaccurate supervision methods often rely on theoretical approximations (e.g., approximate upper and lower bounds), leading to poor model robustness. For instance, as shown in Fig. 2, even with a model size of 10M, EleGANt [40] still frequently produces unsatisfactory results. The above challenges become more pronounced as the model size continues to decrease. The reduced model scale makes the model more sensitive to the accuracy of supervision while also requiring the abandonment of various auxiliary alignment techniques.
@@
To address the above challenges, we propose a novel learning framework named Data Amplify Learning enabling stable and robust training supervision even with a limited number of image annotations. In the heart of DAL is a meticulously designed Diffusion-based Data Amplifier, responsible for amplifying the limited labeled image pairs (typically 5 pairs) into a larger dataset suitable for model training. This process allows DAL to transform the inaccurate supervision into pixel-to-pixel learning (e.g., L1), reducing the optimization difficulty and enabling accurate gradient propagation. Naively using mainstream diffusion-based control methods [41, 35] as the Data Amplifier can lead to subpar outcomes, including identity mismatching, over-smoothing details, and inaccurate makeup results. To alleviate above challenges, two ingredients of the DDA are proposed: (1) Instead of directly generating makeup results, we design a Residual Diffusion Model (RDM) for the Data Amplifier. RDM consists of two branches: one is responsible for face reconstruction, and the other is dedicated to applying makeup. By utilizing the residual between the reconstructed face and the input face, we can transfer this residual to the makeup branch, resulting in high-fidelity and sharpened makeup results. (2) Unlike direct image injection [41], we propose a Fine-Grained Makeup Module (FGMM) that couples fine-grained makeup with facial identity representation to control makeup process, which effectively avoids the loss of facial identity. Additionally, we introduce semantic region labels for makeup, further refining the control of makeup application, rendering a precise makeup control generation. DAL greatly relaxes the optimization methods, allowing us to abandon face prompts and over-parameterization methods, thereby focusing on network design for mobile computing. Benefiting from this, a single TinyBeauty model is designed without reliance on external face prompts. TinyBeauty exceeds prior makeup transfer performance using only 14 convolution layers, enabling efficient deployment on mobile devices. Furthermore, to produce more complete makeup components compared to prior works, we utilize an edge operator to constrain the learning of eyeliner, enhancing the realism and fidelity of generated makeup styles.
@@ 2.1 Diffusion-based Image-to-Image Translation
Compared to previously widely used GAN-based [12, 1, 2, 4, 3, 26] models , recent diffusion models [34, 9, 31] have demonstrated superior performance in high-fidelity image generation tasks, especially when few-shot images are available. Typically, conditions are introduced to steer the image generation process, generally in two forms: text-based and image-based guidance. DALL·E2 [24] and Imagen [32] apply text encoders from pre-trained language models [27, 8] for image generation guidance. Textual Inversion[11] and DreamBooth[30] learn special tokens from example images. However, solely using text for conditional guidance is indirect and inaccurate, making it difficult to ensure stable, consistently generated images. To strengthen conditional constraints, SD Image Variations2 [33] and Stable unCLIP3 [36] directly fine-tune text-conditioned diffusion models on image embeddings. IP-Adapter [41] and InstantID [38] use a combined image and text prompt for data generation conditions to keep the identity of the condition image. While these methods yield commendable outcomes in image stylization, they fail to preserve the facial detail like wrinkles of the original image, thus inadequate in applying facial makeup.
@@
The latency is the inference time on an iPhone 13 device.
@@ 2.2 Facial Makeup
Traditional facial makeup methods [13, 20] rely on facial landmarks to warp predetermined beauty materials on the facial images for makeup application, which is efficient but unrealistic. BeautyGAN [21] introduces a dual GAN approach for makeup transfer with a color histogram loss using the proposed dataset. PSGAN [17] and FAT [37] propose attention mechanisms to address pose/expression changes, significantly increasing model size. SCGAN [6] encodes styles into component-wise codes while EleGANt [40] simplifies this complex optimization problem as L1 loss by generating pseudo ground truth. However, EleGANt still consumes significant computational overhead due to its use of multi-scale attention modules. By abandoning CycleGAN’s structure, BeautyREC [39] achieves a lighter model compared to previous work. Unlike makeup transfer methods, the face beautification proposed in [22] targets many-to-many translation by integrating beauty score prediction.Typically, all these methods incorporate parsing maps and/or landmarks during pre-processing on the unpaired MT dataset, forming cumbersome pipelines.
@@ 3 Method
We propose a novel learning scheme called Data Amplify Learning (DAL) to replace the unstable and inaccurate unpaired learning. DAL leverages a diffusion model to generate previously inaccessible paired data, which is then employed as training material for a compact tiny model. The core of DAL is a Diffusion-based Data Amplifier (DDA) to generate large amounts of paired data using only 5 makeup images as data seeds. (Sec. 3.1). Benefit from the DDA-amplified data, a single TinyBeauty Model is designed to replace the previous cumbersome makeup pipeline, facilitating the integration of the makeup model into mobile devices. (Sec. 3.2).
@@ 3.1 Diffusion-based Data Amplifier
$L_{simple}=\mathbb{E}_{\mathbf{z}_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},t}||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c},t)||^{2},$ (1)
@@
$y=\mathcal{F}_{fine}(x,\mathbf{c}),$ (2)
@@
Residual Diffusion Models To address the issue of facial features like wrinkles being smoothed out during the denoising step of diffusion-based portrait image generation, we introduce a novel framework that utilizes parallel diffusion branches during the inference phase—the conditioned and unconditioned branches. The conditioned branch operates with content condition $\mathbf{c}_{con}$ and style condition $\mathbf{c}_{sty}$ to produce a smooth makeup image, while unconditioned branch proceeds without any conditions to output a smooth non-makeup image.
@@
Fine-Grained Makeup Module The Fine-Grained Makeup Module comprises the Identity Preservation Block (IPB), which safeguards individual facial identity, the Style Preservation Block (SPB) for precise makeup style management, and a latent face mask that guides region-specific makeup application and facilitates the flawless integration of diverse makeup styles.
@@
Mask Guidance. To accurately define the facial areas affected by the makeup style, we divide the feature space into three distinct regions under the guidance of resized facial masks: $M_{face}$, $M_{lips}$, and $M_{eyes}$, as shown in Fig. 4. For single makeup style learning, we define the influenced area $M_{changed}$ as $M_{changed}=M_{face}+M_{lips}+M_{eyes}$. In training process, we only focus the learning on $M_{changed}$ regions to effectively capture makeup style. This is achieved by computing the loss $L_{simple}$ only over the features within the $M_{changed}$ area:
$L_{simple}^{M}=\mathbb{E}_{\mathbf{z}_{0},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},t}||(\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c},t))*M_{changed}||^{2}.$ (4)
@@ 3.2 TinyBeauty Model
$\mathcal{L}_{s}=||\mathcal{S}(y)-\mathcal{S}(y^{\prime})||_{2}^{2}*M_{eyes},$ (5)
@@
$\mathcal{L}_{rec}=||y-y^{\prime}||_{1}.$ (6)
@@ 4.1 Datasets and Evaluation Metrics
Data Seeds and Evaluation Dataset. We randomly select 5 images without makeup from the Flickr-Faces-HQ (FFHQ) [18] dataset as our DAL’s data seeds, i.e. training data. We then manually annotate these data seeds with five predefined makeup styles using MEITU  [^1] https://www.meitu.com/, along with the makeup-only images as the input of SPB. To enable a comprehensive and fair evaluation, 100 images from the FFHQ dataset [18] and 100 images from the MT dataset [21] are additionally annotated as test data when computing PSNR, FID, and LPIPS.
@@
Evaluation Metrics. To evaluate TinyBeauty quantitatively, we employ widely recognized metrics to evaluate the similarity, diversity, and realism of the images generated. We use the Peak Signal-to-Noise Ratio (PSNR) to measure the similarity between the generated data and ground truth (GT) data. A higher PSNR indicates a closer approximation to the GT. Additionally, we report the Fréchet Inception Distances (FID) and Learned Perceptual Image Patch Similarity (LPIPS) by comparing 100 generated images to 100 GT images. FID and LPIPS metrics evaluate the diversity and realism of the generated samples, with lower scores indicating a higher similarity to the GT data distribution.
@@ 4.2 Implementation Details
Diffusion-based Data Amplifier (DDA). Our diffusion fine-tuning experiment use SDv1.5 [25] as the base model. The image encoder utilized is OpenCLIP ViT-H/14 [10], which is pre-trained by IP-Adapter [41]. We set the length of both style token and identity token to 32 to balance the strength of style and identity. LoRA [15] in conjunction with U-net is employed to fine-tune the SD model with a learning rate of 1e-4 for 500 epochs. Notably, the five distinct makeup styles are concurrently trained within a unified model, demonstrating the model’s capacity for multi-style learning in a single training session. The training process is executed on an NVIDIA V100 GPU, taking about 50 mins to fine-tune five makeup styles. We utilize the FaRL [44] to generate facial masks and resize them to a resolution of $64\times 64$ to guide the network’s training in the latent space.
TinyBeauty Model. The TinyBeauty model is trained using a dataset consisting of 4000 diffusion-generated images for 50 epochs. The training processes utilized a learning rate of $42e-$ 4 and the Adam [19] optimizer. The entire training procedure is executed on an NVIDIA V100 GPU, requiring approximately 12 hours. It is important to mention that NO specific face operations are conducted on the input image during the inference phase, which makes our makeup process fast and hardware-friendly.
@@ 4.3 Comparison
To showcase the quality of the data generated using our DDA, we compare it with the text-conditioned SD model, ControlNet [43], IP-Adapter [41], and InstantID [38]. We train both previous methods in the same data as ours, with the result shown in Fig. 5. The results reveal two key insights. Firstly, the subtleties of makeup styles are not easily captured by textual descriptions alone. Consequently, it is challenging to precisely manipulate attributes, particularly color, in the synthesized images. Our DDA addresses this by directly encoding makeup images, thus producing results with consistent color and makeup content. Secondly, IP-Adapter falls short of meeting our stringent criteria for identity retention in the facial makeup task. Our DDA, however, effectively resolves this issue, generating high-quality makeup images by synergistically leveraging mask guidance, IPB, and RDM. Additionally, Fig. 1 illustrates the DDA’s capability of fusing various makeup styles. Our DDA adeptly blends eye, lip, and blush makeup to create a myriad of diverse and intricate makeup applications.
@@
Table 2: Parameters of TinyBeauty and competing methods. (+) means the method uses facial pre-processing including face parsing, and (-) means the method only has a single model.
@@
Quantitative Comparison. We conduct a comprehensive evaluation by comparing TinyBeauty not only with the pre-trained models provided by earlier studies [21, 39, 17, 6, 40], but also by retraining the two most advanced models [40, 39] from those studies using our DAL. The comparative analysis, detailed in Tab. 1, reveals TinyBeauty’s performance over the pre-established and newly trained models in terms of PSNR, FID, and LPIPS across both the FFHQ and MT datasets. Remarkably, despite its significantly smaller network size, TinyBeauty achieves comparable results to EleGANt in the same training set. The success lies in the high-quality paired data we generate, which enables even small networks to learn and apply stable makeup styles effectively. Moreover, our specialized design incorporating residual learning not only diminishes noise interference but also enhances TinyBeauty’s performance, even surpassing larger predecessors. Notably, the MT dataset is not utilized during training, yet TinyBeauty yields a PSNR improvement of 1.49dB over prior methods on MT.
Model Efficiency Evaluation. As shown in Tab. 3, we measure FLOPs, parameter size, and inference time of each model. The models are packaged with CoreML [5] and tested on an iPhone13 without including face pre-processing time. We also test a face parsing model [42] on iPhone13 for reference, commonly utilized by other approaches but not integrated within our TinyBeauty method. Our model achieves an inference time of only 2.18ms for a 256x256 image, which is 13×13$\times$13 × faster than the fastest competing method on iPhone13, i.e., BeautyGAN [21]. Specifically, TinyBeauty is approximately 6 times faster than face parsing on iPhone hardware, which is used as a pre-processing step in previous makeup transfer pipelines [39, 6, 40, 17]. In contrast, our TinyBeauty only utilizes a single model by simplifying the optimization problem facilitated by amplified paired data. These results highlight TinyBeauty’s efficiency and mobile deployment readiness, achieved by implementing our DAL scheme.
@@
User Study. We conduct a subjective evaluation on the FFHQ and MT test sets. Workers are shown makeup results from TinyBeauty and competing methods for raw images in random order. A total of 100 validated workers aged 19-48 from different cultural backgrounds participate in the fair and randomized evaluation. The average rankings are summarized in Tab. 3.